# Unified description of viscous, viscoelastic, or elastic thin active films on substrates

### v1

@@ I Introduction
Active fluids are characterized by their capability of terminal flow. Releasing stresses after longer times of displacement from the original positions leads to persistent finite displacements of the volume elements. Restoring forces that take the volume elements the full way back to their initial positions do not exist. Spontaneously emergent flows in active fluids are directly induced by the enclosed active agents. Primary examples include active nematics [6] and suspensions of microswimmers [3, 4] such as bacteria. They exhibit a rich variety of dynamic spatiotemporal patterns [7, 8, 9]. Apart from complex vortex and swirling states, such as active turbulence [10, 11, 12, 13], one of the most important phenomena is flocking or swarming [14, 15]. These states of polar orientational order and directed collective motion are already found for so-called ‘dry’ active matter. They are described by the seminal Vicsek model [16] and its continuum counterpart, the Toner–Tu theory [17, 18].
@@
Many living biological media do not fit neatly into only one of these two categories of either active viscous fluids or active elastic solids. So far, a number of studies on active fluids have extended considerations on collective motion to more complex rheology. Examples comprise viscoelasticity [32, 33, 34], shear thickening [35], or shear thinning [36]. Several synthetic and many biological materials exhibit both viscous and elastic properties at the same time [37]. For instance, biofilms feature complex rheological behavior [38]. Bacterial biofilms are formed when bacteria excrete extracellular polymeric substances so that they are enclosed by the polymeric material [5, 39]. Depending on the stage of formation of the biofilm, the location within it, and the time scales considered, the material may rather resemble an active fluid or an active solid, but generally show a viscoelastic stage in between. These active biological media thus develop their character from more fluid-like to more solid-like behavior over time or over space.
@@
As a consequence, there is a need for a corresponding unified framework that is capable of characterizing, simply by variations of parameter values, a full range of viscoelastic behavior, from the limit of active fluid- to the limit of active solid-like. Numerous existing continuum theories focus on these two limits of active fluids with simple [18, 40, 12] or complex rheology [32, 33, 34, 35] and active solids [41, 24, 30]. A comprehensive description that can be continuously tuned from an active viscous fluid to an active elastic solid is to be presented.
The aim of our work is to derive such a unified framework. For this purpose, we start from a microscopic picture of active objects self-propelling on a substrate and enclosed by a viscoelastic medium. The underlying microscopic equations are coarse-grained to a continuum description. By variation of a single parameter, our resulting equations can be tuned continuously from the limit of active fluid-like to the limit of active solid-like behavior, covering a broad range of viscoelasticity in between. Additionally, we explore spatially uniform analytical solutions of the derived equations. In this way, we identify a disordered isotropic state for low activity. An orientationally ordered, polar state emerges at elevated activity but weak elasticity. Moreover, a dynamic state of rotating polar orientational order is found for both stronger activity and elasticity. Analytical calculations show how the transition from stationary to rotating states changes from supercritical to subcritical with increasing elasticity. Our analytical results apply to systems of small size under periodic boundary conditions, where the solutions remain spatially uniform.
@@ II Microscopic picture
$\frac{d\mathbf{X}_{n}}{dt}$ $=\mathbf{v}+\sqrt{2D_{\mathrm{tr}}}\boldsymbol{\xi}_{n}\,,$ (1)
@@
$\frac{d\mathbf{N}_{n}}{dt}$ $=\boldsymbol{\Pi}(\mathbf{N}_{n})\cdot\big{[}(\boldsymbol{\Omega}+\tilde{\kappa}\boldsymbol{\Sigma})\cdot\mathbf{N}_{n}+\gamma_{\mathrm{a}}\mathbf{v}+\sqrt{D_{\mathrm{r}}}\boldsymbol{\eta}_{n}\big{]}\,.$ (2)
@@
In these equations, $\mathbf{v}$, $\boldsymbol{\Omega}$, and $\boldsymbol{\Sigma}$ are fields that depend on the position $\mathbf{x}$ within the system and time $t$. $\mathbf{v}$ denotes the locally averaged total velocity field of the suspension at position $\mathbf{x}$ and time $t$, which will be discussed in detail in Sec. IV.
@@
The first equation describes advective transport of the active objects with velocity $\mathbf{v}$ and additionally includes a Gaussian white noise $\boldsymbol{\xi}_{n}$ with a translational diffusion coefficient $D_{\mathrm{tr}}$. Activity enters via the forces that the self-propelled objects exert on the surrounding medium to set it and themselves into motion, see below.
@@
Equation (2) quantifies the dynamics of the unit vector $\mathbf{N}_{n}$ that encodes the current orientation of object $n$. Within the square brackets, the first contribution describes reorientations by the local vorticity of the flow field, $\boldsymbol{\Omega}=[(\nabla\mathbf{v})^{\top}-(\nabla\mathbf{v})]/2$ [42]. Similar effects result from distortional gradients of the flow as included by the deformation rate tensor $\boldsymbol{\Sigma}=[(\nabla\mathbf{v})^{\top}+(\nabla\mathbf{v})]/2$ [42]. The parameter $\tilde{\kappa}$ denotes the tumbling parameter and is a function of the ratio $r$ of major to minor axis of the body of the active object. More precisely, $\tilde{\kappa}=(r^{2}-1)/(r^{2}+1)$. Thus, $\tilde{\kappa}=0$ for spherical and $\tilde{\kappa}>0$ for elongated objects [42]. Next, the term $\gamma\mathbf{v}_{a}$ describes alignment with the direction of the overall flow. For this effect to emerge, the self-propelled units need to be able to “hold on” to something to withstand pure advection by the flow, such as a supporting substrate in thin active films [43, 44, 45]. Assuming that the active agents exhibit anisotropic friction with respect to the enclosing medium and are capable of reorienting in flows, movement relative to the medium induces rotations [46]. Specifically, active agents experiencing higher friction at the rear than at the front reorient into the direction of the flow, and vice versa. Our description takes into account this “weathervane” effect already on the microscopic level. We remark that a similar term is included in various discretized models of active solids [47, 48, 26, 49]. Equation (2) adds a Gaussian white noise term $\boldsymbol{\eta}_{n}$ to the current orientation with rotational diffusion constant $D_{\mathrm{r}}$. Finally, the overall equation is multiplied by the projector $\boldsymbol{\Pi}(\mathbf{N}_{n})=\mathbf{I}-\mathbf{N}_{n}\mathbf{N}_{n}$, where $\mathbf{I}$ refers to the unit matrix. This projection maintains the nature of $\mathbf{N}_{n}$ as a unit vector.
@@ III Polar orientational order parameter
$-\nabla_{\mathbf{n}}\cdot\big{\{}\big{[}\boldsymbol{\Omega}\cdot\mathbf{n}+\tilde{\kappa}\boldsymbol{\Pi}(\mathbf{n})\cdot\boldsymbol{\Sigma}\cdot\mathbf{n}$
$\qquad+\gamma_{\mathrm{a}}\boldsymbol{\Pi}(\mathbf{n})\cdot\mathbf{v}-D_{\mathrm{r}}\mathbf{n}\big{]}\mathcal{P}(\mathbf{x},\mathbf{n},t)\big{\}}$
$+\frac{1}{2}D_{\mathrm{r}}\nabla_{\mathbf{n}}\nabla_{\mathbf{n}}:[\boldsymbol{\Pi}(\mathbf{n})\mathcal{P}(\mathbf{x},\mathbf{n},t)]\,.$ (4)
Here, the first two terms describe translational drift and translational diffusion, respectively. The next two terms emerge due to rotational drift and reorientation under the influence of the flow $\mathbf{v}$ of the surrounding medium. Rotational diffusion causes the contribution by the last term. Again, we use the abbreviation $\boldsymbol{\Pi}(\mathbf{n})=\mathbf{I}-\mathbf{n}\mathbf{n}$.
@@
$\partial_{t}\mathbf{P}$ $=$ ${}-\nabla\cdot\langle\mathbf{v}\mathbf{n}\rangle+D_{\mathrm{tr}}\nabla^{2}\mathbf{P}-D_{\mathrm{r}}\mathbf{P}+\langle\boldsymbol{\Omega}\cdot\mathbf{n}\rangle$ (10)
${}+\tilde{\kappa}\langle(\mathbf{I}-\mathbf{n}\mathbf{n})\cdot\boldsymbol{\Sigma}\cdot\mathbf{n}\rangle+\gamma_{\mathrm{a}}\langle(\mathbf{I}-\mathbf{n}\mathbf{n})\cdot\mathbf{v}\rangle\,.\quad$
@@
$\partial_{t}\mathbf{P}=$ $-\mathbf{v}\cdot\nabla\mathbf{P}+D_{\mathrm{tr}}\nabla^{2}\mathbf{P}-D_{\mathrm{r}}\mathbf{P}+\boldsymbol{\Omega}\cdot\mathbf{P}$ (11)
$+\tilde{\kappa}\boldsymbol{\Sigma}\cdot\mathbf{P}-\tilde{\kappa}\boldsymbol{\Sigma}:\langle\mathbf{n}\mathbf{n}\mathbf{n}\rangle+\gamma_{\mathrm{a}}\mathbf{v}-\gamma_{\mathrm{a}}\mathbf{v}\cdot\langle\mathbf{n}\mathbf{n}\rangle\,.$
@@
The third moment is treated via the so-called Hand-closure [54, 53], see Ref. 12 for a more detailed discussion. It implies
@@
$\partial_{t}\mathbf{P}=$ $-\mathbf{v}\cdot\nabla\mathbf{P}+\boldsymbol{\Omega}\cdot\mathbf{P}+\kappa\boldsymbol{\Sigma}\cdot\mathbf{P}+D_{\mathrm{tr}}\nabla^{2}\mathbf{P}-D_{\mathrm{r}}\mathbf{P}$
@@ IV Hydrodynamics and elasticity
Having addressed the orientational order parameter fields, we next turn to the coupling between Eq. (III) and the overall velocity field $\mathbf{v}$ of the suspension. Under low-Reynolds number conditions and when exposed to linear friction with a supporting substrate, we evaluate the equation
@@
This amended Stokes equation describes the balance of multiple forces.
First, we assume the material to be incompressible, $\nabla\cdot\mathbf{v}=0$. As a result, the pressure field $p$ acts as a Lagrange multiplier enforcing incompressibility. Second, viscous forces are quantified by shear viscosity $\eta$. To take into account elastic forces, we introduce the displacement field $\mathbf{u}$. It includes the reversible elastic displacements at time $t$ stored for the material elements currently located at positions $\mathbf{x}$. Thus, if all stresses were released at time $t$, the material elements would tend to return to positions $\mathbf{x}-\mathbf{u}(\mathbf{x},t)$. We discuss the equation governing the dynamics of $\mathbf{u}$ below and in Appendix A. The shear modulus $\mu$ is a measure for the mechanical stiffness of the material, that is, its resistance against elastic deformations.
Since we consider thin films of active suspensions on substrates, additional frictional forces linear in overall velocity $\mathbf{v}$ are taken into account. $\nu_{\mathrm{v}}$ is the corresponding friction coefficient. We further allow for elastic anchoring to the supporting substrate, which leads to elastic restoring forces. They tend to drive the material elements back to elastically memorized positions as quantified by the displacement field $\mathbf{u}$. We consider corresponding restoring forces to be linear in $\mathbf{u}$. Their strength is set by the parameter $\nu_{\mathrm{d}}$.
@@
Here, we have introduced the strength of active forcing, $\nu_{\mathrm{p}}=cf_{a}$. This kind of contribution corresponds to propulsive active forces often employed when modeling active solids [24, 20, 55, 27].
@@
Overall, we aim to derive a theoretical description of active materials that covers both viscous fluids and elastic solids, as well as a broad range of viscoelastic materials in between. In Eq. (18), the displacement field $\mathbf{u}$ includes the aspects of elasticity. It can be viewed as a memory field. Through $\mathbf{u}$, the material element at each position $\mathbf{x}$ memorizes a position $\mathbf{x}-\mathbf{u}(\mathbf{x},t)$, to where it would return if all stresses were instantly relaxed at time $t$. Generally, in viscoelastic materials, this memory associated with the displacement field may fade away over time. In other words, such materials are not purely elastic. They flow, which implies that $\mathbf{u}$ relaxes over time. In the simplest case,
@@
In this equation, the first contribution $\mathbf{v}$ on the right-hand side describes how displacements $\mathbf{u}$ arise due to transport of the material. Contrarily, the second term quantifies how the memory associated with the displacement field is lost over time. The decay rate of the memory is set by the inverse relaxation time $\tau_{\mathrm{d}}$. Furthermore, the left-hand side of Eq. (20) additionally includes convective transport of $\mathbf{u}$ by material flow. This contribution represents a first step beyond linear elasticity, particularly in combination with Eq. (18). Such transport becomes important when considering viscoelastic fluids that are able to support persistent flow.
The derivation of Eq. (20) together with associated assumptions are summarized in Appendix A. There, the origin of the field $q$ is detailed as well. Its contribution via $-\nabla q$ emerges when consistently maintaining incompressibility $\nabla\cdot\mathbf{u}=0$ of the system while linearizing strains under the action of the nonlinear convective term $\mathbf{v}\cdot\nabla\mathbf{u}$.
Varying the relaxation time $\tau_{\mathrm{d}}$ of the elastic memory, we can continuously tune the behavior of the material from the limit of purely fluid-like to the opposite limit of purely elastic, solid-like. More precisely, the limit of a viscous fluid is obtained for instant relaxation $\tau_{\mathrm{d}}\to 0$. The latter implies absence of elastic memory. In this case, Eq. (20) simply yields $\mathbf{u}=\mathbf{0}$ and Eq. (18) reduces to
$\mathbf{0}=-\nabla p+\eta\nabla^{2}\mathbf{v}-\nu_{\mathrm{v}}\mathbf{v}+\nu_{\mathrm{p}}\mathbf{P}\,.$ (21)
Contrarily, the limit $\tau_{\mathrm{d}}\to\infty$ of diverging relaxation time describes a non-decaying elastic memory. In other words, the behavior is perfectly elastic. Then, all displacements are completely reversible. Still, in this solid-like case, damping by viscous and frictional forces affects the dynamics.

### v2

@@ I Introduction
Active fluids are characterized by their capability of terminal flow. Releasing stresses after longer times of displacement from the original positions leads to persistent finite displacements of the volume elements. Restoring forces that take the volume elements the full way back to their initial positions do not exist. Spontaneously emergent flows in active fluids are directly induced by the enclosed active agents. Primary examples include active nematics [6] and suspensions of microswimmers [3, 4] such as bacteria. They exhibit a rich variety of dynamic spatiotemporal patterns [7, 8, 9, 10]. Apart from complex vortex and swirling states, such as active turbulence [11, 12, 13, 14], one of the most important phenomena is flocking or swarming [15, 16]. These states of polar orientational order and directed collective motion are already found for so-called ‘dry’ active matter. They are described by the seminal Vicsek model [17] and its continuum counterpart, the Toner–Tu theory [18, 19].
@@
Numerous existing continuum theories focus on these two limits of active fluids with simple [19, 33, 13] or complex rheology [34, 35, 36, 37, 38] and active solids [39, 25, 31]. Yet, many living biological media do not fit neatly into only one of these two categories of either active viscous fluids or active elastic solids. So far, a number of studies on active fluids have extended considerations on collective motion to more complex rheology. Examples comprise viscoelasticity [34, 35, 36, 40] or shear thickening and thinning [41, 37, 38]. In particular, the interplay between viscoelasticity and activity in fluid-like systems, concerning their long-term behavior, has been explored in the past to some extent, focusing, for example, on nematic systems [34, 35], multicomponent gels [42], and the motion of probe particles [43].
Correspondingly, several synthetic and many biological materials exhibit both viscous and elastic properties at the same time [44]. For instance, biofilms feature complex rheological behavior [45]. Bacterial biofilms are formed when bacteria excrete extracellular polymeric substances so that they are enclosed by the polymeric material [5, 46]. Depending on the stage of formation of the biofilm, the location within it, and the time scales considered, the material may rather resemble an active fluid or an active solid, but generally show a viscoelastic stage in between. These active biological media thus develop their character from more fluid-like to more solid-like behavior over time or over space.
@@
As a consequence, there is a need for a corresponding unified framework that is capable of characterizing, simply by variations of parameter values, a full range of viscoelastic behavior, from the limit of active fluid- to the limit of active solid-like. A comprehensive description combining activity, entering via the self-propulsion of active agents, and viscoelastic material behavior that can be continuously tuned from an active viscous fluid to an active elastic solid is to be presented.
The aim of our work is to derive such a unified framework. For this purpose, we start from a microscopic picture of active objects self-propelling on a substrate and enclosed by a viscoelastic medium. The underlying microscopic equations are coarse-grained to a continuum description. By variation of a single parameter, our resulting equations can be tuned continuously from the limit of active fluid-like to the limit of active solid-like behavior, covering a broad range of viscoelasticity in between. Additionally, we explore spatially uniform analytical solutions of the derived equations. In this way, we identify a disordered isotropic state for low activity. An orientationally ordered, polar state emerges at elevated activity but weak elasticity. Moreover, a dynamic state of rotating polar orientational order is found for both stronger activity and elasticity. Analytical calculations show how the transition from stationary to rotating states changes from supercritical to subcritical with increasing elasticity. Our analytical results apply to systems of small size under periodic boundary conditions, where the solutions remain spatially uniform. The more general case is referred to numerically in a companion article [47] and demonstrates spatial decorrelation for larger systems.
@@ II Microscopic picture
$\frac{d\mathbf{X}_{n}}{dt}$ $=\mathbf{v}+\sqrt{2D_{\mathrm{tr}}}\bm{\xi}_{n}\,,$ (1)
@@
$\frac{d\mathbf{N}_{n}}{dt}$ $=\bm{\Pi}(\mathbf{N}_{n})\cdot\big{[}(\bm{\Omega}+\tilde{\kappa}\bm{\Sigma})\cdot\mathbf{N}_{n}+\gamma_{\mathrm{a}}\mathbf{v}+\sqrt{D_{\mathrm{r}}}\bm{\eta}_{n}\big{]}\,.$ (2)
@@
In these equations, $\mathbf{v}$, $\bm{\Omega}$, and $\bm{\Sigma}$ are fields that depend on the position $\mathbf{x}$ within the system and time $t$. $\mathbf{v}$ denotes the locally averaged total velocity field of the suspension at position $\mathbf{x}$ and time $t$, which will be discussed in detail in Sec. IV.
@@
The first equation describes advective transport of the active objects with velocity $\mathbf{v}$ and additionally includes a Gaussian white noise $\bm{\xi}_{n}$ with a translational diffusion coefficient $D_{\mathrm{tr}}$. Activity enters via the forces that the self-propelled objects exert on the surrounding medium to set it and themselves into motion, see below.
@@
Equation (2) quantifies the dynamics of the unit vector $\mathbf{N}_{n}$ that encodes the current orientation of object $n$. Within the square brackets, the first contribution describes reorientations by the local vorticity of the flow field, $\bm{\Omega}=[(\nabla\mathbf{v})^{\top}-(\nabla\mathbf{v})]/2$ [48]. Similar effects result from distortional gradients of the flow as included by the deformation rate tensor $\bm{\Sigma}=[(\nabla\mathbf{v})^{\top}+(\nabla\mathbf{v})]/2$ [48]. The parameter $\tilde{\kappa}$ denotes the tumbling parameter and is a function of the ratio $r$ of major to minor axis of the body of the active object. More precisely, $\tilde{\kappa}=(r^{2}-1)/(r^{2}+1)$. Thus, $\tilde{\kappa}=0$ for spherical and $\tilde{\kappa}>0$ for elongated objects [48]. Next, the term $\gamma\mathbf{v}_{a}$ describes alignment with the direction of the overall flow. For this effect to emerge, the self-propelled units need to be able to “hold on” to something to withstand pure advection by the flow, such as a supporting substrate in thin active films [49, 50, 51]. Assuming that the active agents exhibit anisotropic friction with respect to the enclosing medium and are capable of reorienting in flows, movement relative to the medium induces rotations [52]. Specifically, active agents experiencing higher friction at the rear than at the front reorient into the direction of the flow, and vice versa. Our description takes into account this “weathervane” effect already on the microscopic level. We remark that a similar term is included in various discretized models of active solids [53, 54, 27, 55]. Equation (2) adds a Gaussian white noise term $\bm{\eta}_{n}$ to the current orientation with rotational diffusion constant $D_{\mathrm{r}}$. Finally, the overall equation is multiplied by the projector $\bm{\Pi}(\mathbf{N}_{n})=\mathbf{I}-\mathbf{N}_{n}\mathbf{N}_{n}$, where $\mathbf{I}$ refers to the unit matrix. This projection maintains the nature of $\mathbf{N}_{n}$ as a unit vector.
Our study focuses on systems in which the interactions between the self-propelled objects are primarily mediated by the continuous medium in between. This seems justified for low to intermediate concentrations of active units, that is, when these units rarely approach closely to each other. For higher concentrations, additional interactions, such as steric repulsion, should be considered as well.
@@ III Polar orientational order parameter
$-\nabla_{\mathbf{n}}\cdot\big{\{}\big{[}\bm{\Omega}\cdot\mathbf{n}+\tilde{\kappa}\bm{\Pi}(\mathbf{n})\cdot\bm{\Sigma}\cdot\mathbf{n}$
$\qquad+\gamma_{\mathrm{a}}\bm{\Pi}(\mathbf{n})\cdot\mathbf{v}-D_{\mathrm{r}}\mathbf{n}\big{]}\mathcal{P}(\mathbf{x},\mathbf{n},t)\big{\}}$
$+\frac{1}{2}D_{\mathrm{r}}\nabla_{\mathbf{n}}\nabla_{\mathbf{n}}:[\bm{\Pi}(\mathbf{n})\mathcal{P}(\mathbf{x},\mathbf{n},t)]\,.$ (4)
Here, the first two terms describe translational drift and translational diffusion, respectively. The next term emerges due to rotational drift under the influence of the flow $\mathbf{v}$ of the surrounding medium. Rotational diffusion causes the contribution by the last term. Again, we use the abbreviation $\bm{\Pi}(\mathbf{n})=\mathbf{I}-\mathbf{n}\mathbf{n}$.
@@
$\partial_{t}\mathbf{P}$ $=$ ${}-\nabla\cdot\langle\mathbf{v}\mathbf{n}\rangle+D_{\mathrm{tr}}\nabla^{2}\mathbf{P}-D_{\mathrm{r}}\mathbf{P}+\langle\bm{\Omega}\cdot\mathbf{n}\rangle$ (10)
${}+\tilde{\kappa}\langle(\mathbf{I}-\mathbf{n}\mathbf{n})\cdot\bm{\Sigma}\cdot\mathbf{n}\rangle+\gamma_{\mathrm{a}}\langle(\mathbf{I}-\mathbf{n}\mathbf{n})\cdot\mathbf{v}\rangle\,.\quad$
@@
$\partial_{t}\mathbf{P}=$ $-\mathbf{v}\cdot\nabla\mathbf{P}+D_{\mathrm{tr}}\nabla^{2}\mathbf{P}-D_{\mathrm{r}}\mathbf{P}+\bm{\Omega}\cdot\mathbf{P}$ (11)
$+\tilde{\kappa}\bm{\Sigma}\cdot\mathbf{P}-\tilde{\kappa}\bm{\Sigma}:\langle\mathbf{n}\mathbf{n}\mathbf{n}\rangle+\gamma_{\mathrm{a}}\mathbf{v}-\gamma_{\mathrm{a}}\mathbf{v}\cdot\langle\mathbf{n}\mathbf{n}\rangle\,.$
@@
The third moment is treated via the so-called Hand-closure [60, 59], see Ref. 13 for a more detailed discussion. It implies
@@
$\partial_{t}\mathbf{P}=$ $-\mathbf{v}\cdot\nabla\mathbf{P}+\bm{\Omega}\cdot\mathbf{P}+\kappa\bm{\Sigma}\cdot\mathbf{P}+D_{\mathrm{tr}}\nabla^{2}\mathbf{P}-D_{\mathrm{r}}\mathbf{P}$
@@ IV Hydrodynamics and elasticity
Having addressed the orientational order parameter fields, we next turn to the coupling between Eq. (III) and the overall velocity field $\mathbf{v}$ of the suspension as well as the degrees of freedom associated with elasticity. The systematic path of introducing further degrees of freedom, particularly those associated with spontaneous symmetry breaking, into the theory of hydrodynamics was outlined more than half a century ago [61]. Specifically, this concerns broken translational symmetry in elastic solids, associated with an additional field of displacements $\mathbf{u}$. Its extension to viscoelastic systems was systematically derived within this framework more than a quarter of a century ago [62]. Here, we follow this path of generalized hydrodynamics and employ it to analyze the properties of active viscoelastic suspensions on substrates.
It is the Euler point of view that is associated with generalized hydrodynamics. This perspective allows for the interpretation within the hydrodynamics framework of what is usually referred to as displacement field $\mathbf{u}$ in elasticity theory. Introducing elasticity into generalized hydrodynamics, $\mathbf{u}$ can be viewed as the field of currently and locally “memorized” elastic displacements in the system [63].
Usually, one starts the list of hydrodynamic equations by the continuity equation, that is, the dynamic equation for the density. We consider the active viscoelastic material as incompressible with approximately constant density. For biological and particularly aqueous systems, this frequently represents a reasonable assumption. Thus, the hydrodynamic continuity equation simplifies to $\nabla\cdot\mathbf{v}=0$.
Next, we turn to the dynamic equation for the flow velocity itself. For simple liquids, it is given by the Navier-Stokes equation. We here consider low-Reynolds number conditions and overdamped dynamics. Additional linear friction with a supporting substrate is included. Therefore, the inertial terms are neglected, and we evaluate an amended Stokes equation,
@@
It describes the balance of multiple forces, which we explain in the following.
On the right-hand side, for incompressible systems, the pressure field $p$ effectively acts as a Lagrange multiplier enforcing incompressibility $\nabla\cdot\mathbf{v}=0$. The second term introduces the viscous forces quantified by the shear viscosity $\eta$, as usually in the Stokes equation. This contribution is developed from the divergence of the stress tensor, $\nabla\cdot\bm{\sigma}$, for isotropic, incompressible fluids [64]. Yet, if the system shows additional elastic features, also elastic stresses $\bm{\sigma}^{\mathrm{el}}$ contribute to the overall hydrodynamic stress. For isotropic, incompressible systems $\bm{\sigma}^{\mathrm{el}}=2\mu\mathbf{U}$, where $\mathbf{U}$ denotes the strain tensor. In linear elasticity, it is typically expressed in terms of the elastic displacement field $\mathbf{u}(\mathbf{x},t)$. The latter includes the reversible elastic displacements stored (“memorized”) at time $t$ for the material elements currently located at positions $\mathbf{x}$. Moreover, in linear elasticity theory $\mathbf{U}=\left(\nabla\mathbf{u}+(\nabla\mathbf{u})^{\mathrm{T}}\right)/2$ [65, 66, 67], while incompressible systems satisfy $\nabla\cdot\mathbf{u}=0$ [65]. Together, we recover from $\nabla\cdot\bm{\sigma}^{\mathrm{el}}$ the third term on the right-hand side of Eq. (18). The shear modulus $\mu$ is a measure for the mechanical stiffness of the material, that is, its resistance against elastic deformations.
Since we consider thin films of active suspensions on substrates, additional frictional forces linear in overall velocity $\mathbf{v}$ are taken into account. $\nu_{\mathrm{v}}$ is the corresponding friction coefficient. We further allow for elastic anchoring or pinning to the supporting substrate, which leads to elastic restoring forces. They tend to drive the material elements back to elastically memorized positions as quantified by the field $\mathbf{u}$. We here consider corresponding restoring forces to be linear as well and set their strength by the parameter $\nu_{\mathrm{d}}$.
@@
We introduced the strength of active forcing, $\nu_{\mathrm{p}}=cf_{a}$. This kind of contribution corresponds to propulsive active forces often employed when modeling active solids [25, 21, 68, 28].
@@
In the systematic approach of generalized hydrodynamics, each variable associated with a spontaneously broken symmetry requires and follows a separate dynamic equation as well [61]. Integrating elasticity into generalized hydrodynamics, this spontaneously broken symmetry is reflected by elastic displacements. In Refs. 62, 67, the associated dynamic equation was formulated in terms of the elastic strain tensor $\mathbf{U}$. The correspondence between this tensorial approach and a formulation in terms of a vectorial quantity was pointed out [62], see also Appendix A.
Using again the identification $\mathbf{U}=\left(\nabla\mathbf{u}+(\nabla\mathbf{u})^{\mathrm{T}}\right)/2$, the systematically derived basic dynamic equation resulting within the framework of generalized hydrodnamics for $\mathbf{U}$ [67] can be reduced to the vectorial dynamic equation
@@
In brief, full linearization leads to $\partial_{t}\mathbf{u}=\mathbf{v}-\tau_{\mathrm{d}}^{-1}\mathbf{u}$, which was presented before [63, 71]. In this expression, generation of elastic, memorized displacements $\mathbf{u}$ through motion of material elements $\mathbf{v}$ is obvious. At the same time, in viscoelastic substances internal stress and strain or, correspondingly, memorized elastic displacements, may decay over time. This contribution was outlined in the generalized hydrodynamic theory as well [62, 67]. For example, strained entangled polymeric melts may relax their stress and strain through disentanglement processes [72, 57]. The decay rate of this elastic memory is set by the inverse relaxation time $\tau_{\mathrm{d}}$.
Furthermore, in Eq. (20), we perform one step beyond the completely linearized picture. We additionally keep, during linearization of the nonlinear description of generalized hydrodynamics [62, 67], the term of convective transport $\mathbf{v}\cdot\nabla\mathbf{u}$. Still, the equation remains linear in $\mathbf{u}$. Such transport becomes important when considering viscoelastic fluids that are able to support persistent flow. On the right-hand side of Eq. (20), it requires the additional contribution $-\nabla q$ that emerges when consistently maintaining incompressibility $\nabla\cdot\mathbf{u}=0$ of the system when we keep the description linear in $\mathbf{u}$. Further details on Eq. (20) together with associated assumptions are summarized in Appendix A.
In this context, we review once more the interpretation of the field of elastic displacements $\mathbf{u}$ in the generalized hydrodynamics framework. We recall that the generalized hydrodynamics approach is genuinely field-based, not trajectory-based. In other words, it takes the Euler perspective, instead of the Lagrange point of view [66, 62, 73]. The Euler perspective quantifies the material in its current state at time $t$, at each position $\mathbf{x}$ in space. Accordingly, the field of elastic displacements $\mathbf{u}(\mathbf{x},t)$ does not describe to which position a material element moves from position $\mathbf{x}$. Instead, it quantifies from which position $\mathbf{x}-\mathbf{u}(\mathbf{x},t)$ the material element presently located at position $\mathbf{x}$ has come from. Correspondingly, Ref. 62 in the generalized hydrodynamics framework bases the theory on an initial field $\bm{a}(\mathbf{x},t)=\mathbf{x}-\mathbf{u}(\mathbf{x},t)$. In this sense, $\mathbf{u}$ encodes the memorized positions, from where the material elements have moved to $\mathbf{x}$. To express these relations and emphasize the Euler perspective, we refer to $\mathbf{u}$ as the elastic memory field or field of memorized elastic displacements. Thus, if all stresses were released at time $t$, the material elements would tend to return to positions $\mathbf{x}-\mathbf{u}(\mathbf{x},t)$.
This “retrospective” Euler point of view leads to marked differences of interpretation of the quantities when compared to Lagrange descriptions. Specifically, this becomes apparent in the context of viscoelasticity. Above, we mentioned the relaxation of entangled polymer melts when stretched into a strained state and maintained in this situation. Elastic stress and strain are directly related to each other in linear elasticity for isotropic, incompressible materials via $\bm{\sigma}^{\mathrm{el}}=2\mu\mathbf{U}=\mu\left(\nabla\mathbf{u}+(\nabla\mathbf{u})^{\mathrm{T}}\right)$, see also Appendix A. Relaxation of stress and strain in a viscoelastic system therefore implies decay of $\mathbf{u}$. Naturally, the space positions $\mathbf{x}$ of physical space, where $\mathbf{u}(\mathbf{x},t)$ is defined and evaluated, remain unaffected by this process of relaxation. Thus, it is $\mathbf{x}-\mathbf{u}(\mathbf{x},t)$ that changes. These are the memorized positions, where the material elements would displace back to upon release of stresses. In the Euler perspective, the memorized positions approach the current locations $\mathbf{x}$ during viscoelastic relaxation. The material elements presently located at $\mathbf{x}$ would shift back to positions closer to $\mathbf{x}$ than before relaxation. For a basic graphical illustration we refer to a recent review on the generalized hydrodynamics approach for passive systems [73]. In other words, the system forgets about the initial unstrained state. Its elastic memory, quantified by $\mathbf{u}$, decays so that the system relaxes into a new unstrained state. The rate of decay of this memory is $\tau_{\mathrm{d}}^{-1}$ in Eq. (20).
Varying the relaxation time $\tau_{\mathrm{d}}$ of the elastic memory, we can continuously tune the behavior of the material from the limit of purely fluid-like to the opposite limit of purely elastic, solid-like. More precisely, the limit of a viscous fluid is obtained for instant relaxation $\tau_{\mathrm{d}}\to 0$. Contrarily, the limit $\tau_{\mathrm{d}}\to\infty$ of diverging relaxation time describes a non-decaying elastic memory. In this case, the behavior is perfectly elastic and all displacements are completely reversible. Still, in this solid-like case, damping by viscous and frictional forces affects the dynamics. Indeed, Eq. (20) in its linearized version together with the stress-strain relation $\bm{\sigma}^{\mathrm{el}}=\mu\mathbf{U}$ directly reproduces what is frequently referred to as the basic constitutive linear relations of viscoelasticity. We obtain the Maxwell model for finite $\tau_{\mathrm{d}}$ and $\eta=0$ [67, 73]. For $\tau_{\mathrm{d}}\rightarrow\infty$, we find the Kelvin-Voigt model [73]. This also implies that, by adjusting one single parameter $\tau_{\mathrm{d}}$, we can tune the description from viscous fluid-like, via a broad range of viscoelastic behavior or imperfect elasticity in between, to genuinely elastic systems.